# Inflated hot Jupiters: inferring average atmospheric velocity via Ohmic models coupled with internal dynamo evolution

### v1

@@ 4.3 Estimates of dynamo magnetic fields
where $V$, $r_{i}$ and $r_{o}$ refer to the volume, inner and outer boundary of the convective region, $\mu_{0}$ is the magnetic permeability, $L={\rm min}(r_{o}-r_{i},H_{\rho})$, where $H_{\rho}$ is the density scale height, $H_{T}$ is the temperature scale height, $q_{c}(r)$ is the convective heat flux, and $f_{\rm ohm}$ is the fraction of Ohmic to total dissipation in the dynamo region (which Christensen & Aubert (2006) find numerically to be in the range $\sim 0.3$-$0.7$, depending on the simulation). We evaluate the integral taking $r_{i}$ as the bottom of the envelope (outside the inert core), and $r_{o}=R_{\rm dyn}$, i.e. the radius corresponding to a pressure $p_{\rm dyn}:=10^{6}$ bar, taken as a proxy for the surface of the deep high-conductivity region hosting the dynamo. Note that, strictly speaking, the transition to the metallic hydrogen region corresponds to a $\rho$-$T$ line in the phase diagram, so that $p_{\rm dyn}$ depends slightly on $T$ (or $\rho$). However, since the evaluation of the field is performed via an integral, Eq. (18), a slight change of the upper bound $r_{o}$ implies minor modifications, certainly smaller than the overall intrinsic uncertainties of the scaling law. We checked that, with this recipe, we recover reasonable values of $B_{\rm dyn}\sim 30-50$ G for a Jupiter-like model (negligible irradiation, $M=1\leavevmode\nobreak\ M_{j}$), at $4.5$ Gyr-old. In our code, we use $f_{\rm ohm}\sim 0.5$, in the middle of the typical range $f_{\rm ohm}\sim 0.3-0.8$ of e.g. the geodynamo simulations of Christensen & Aubert (2006) and the anelastic ones by Gastine & Wicht (2021), and on the upper range of cold Jupiter simulations by Elias-López et al. (2025b). In any case, note that the systematic uncertainties in the absolute values of $B_{\rm dyn}(t)$ (which we can overall quantify of a factor $\sim 2$) affect the inferred values of $v_{\rm avg}$, but don’t affect the general trends of the magnetic field evolution. We refer the reader to App. A for a comparison of different versions of the scaling law (Reiners et al., 2009; Reiners & Christensen, 2010), and some words of caution about their applicability to the HJ scenario.
Moreover, we can estimate the average value of the magnetic field intensity at the surface (close to where winds circulate), which we use in eq. (10), by considering that the weights $b_{l}(R_{\rm dyn})$ (see Sect. 4.2.2):
@@
where the last equation considers as a only free parameter the dipolarity at the dynamo surface, $f_{\rm dip}$, for simplicity (and given the much larger source of uncertainties in the scaling law). The set of anelastic dynamo simulations of Elias-López et al. (2025b) found that the dipolarity (calculated over the entire dynamo region, or at its outer surface) is seen to suddenly increase at a few Gyr, switching from a multipolar- to a dipolar-dominated dynamo (see also Yadav et al. 2013 for a discussion of these two branches). Here, for simplicity, we employ $f_{\rm dip}=0.1$, constant in time, in the range of what found by Elias-López et al. (2025b). According to that study, neglecting its time variation might overestimate the surface field at early ages. Using these values of $f_{\rm ohm}$ and $f_{\rm dip}$ allows us to roughly recover Jupiter values of magnetic field at the surface.
@@ 4.4 Numerical details about the heating term
We gradually activate the Ohmic source term $\epsilon_{j}$, eq. (6), with a linear increase in time starting from zero at 5 Myr, to have it fully operative at 20 Myr. The reasons to neglect the Ohmic term at earlier ages are: (i) we only consider HJs with $t>100$ Myr, discarding the few younger ones (see Sect. 2); (ii) planetary formation, migration, and interaction with the disk are still happening, and the Ohmic model doesn’t account for them; (iii) the internal structure and the radius is still sensitive to the (poorly constrained) initial entropy (i.e., initial radius), something irrelevant after $\sim 100$ Myr; (iv) at tless-than-or-similar-$t\lesssim$ Myr, the different versions of the scaling law differ substantially between each other (see App. A), and eq. (18) easily gives unphysical Ohmic efficiencies due to the large heat flux (which in turn is affected by the choice of the initial entropy).
@@
Figure 5: Evolution of the radius $R$ for planets with a given $T_{\rm eq}=1750,2000,2250$ K (from left to right), different masses ($M=0.7,1,2,4$ $M_{j}$ in blue, green, orange, and purple, respectively), and different values of $v_{\rm avg}$, marked by increasingly darker shades: $v_{\rm avg}=\{0,1000,2000,3000\}/[(M/M_{j})(T_{\rm eq}/1500{\rm K})^{6}]$ m/s. In each panel, the grey band indicates the range of observed radii for HJs with $T_{\rm eq}$ around (±$\pm 125$ K) the value assumed in the models, $\bar{R}\pm\sigma_{R}$, as shown by grey background boxes in Fig. 1.
@@
Due to the non-trivial feedback between the background field and the internal heat, the code can show instabilities (for both numerical and physical reasons, see below) for high level of irradiation and/or heat deposition, visible as perturbations in the long-term evolutionary curves, and possibly implying early crashes in the simulations. To reduce the numerical artifacts, we use a reduced time and grid resolution, typically 2x or 4x the default one (i.e., delta_time_coeff and mesh_delta_coeff 0.5 or 0.25 instead on 1). We also implement an exponential decrease of the heat source for $p\lesssim 1$ bar, which helps the convergence in the structure solution and doesn’t affect the results in terms of inflated radii (the fraction of the heat deposited in these outermost thin layers is minimal).
@@
Additionally, the source $\epsilon_{j}$, at a given time $t$, can be actually implemented as the average of the Ohmic heat $<\frac{Q^{n}_{j}}{\rho^{n}}>_{t_{\rm avg}}$ over an interval much smaller than the typical cooling timescales, which are comparable to the age itself (we checked that such a time average doesn’t affect the overall long-term trend of the radius evolution). This simplified multistep-like method, adapted to our case, smooths out the oscillations in the instantaneous Ohmic heating rate, which are due to intermittent convection and dynamo field (see Sect. 5.2) and stabilize the code. In other words, we numerically mitigate the stiffness of the source term. Notice that, in our specific case, the use of the implicit scheme for extra energy sources (with the recently implemented auto differentiation, Jermyn et al. 2023) doesn’t improve the stability due to the intrinsically variable instantaneous heat source, therefore we recurred to the more manual solution described above.
@@ 5.1 Dependence on $T_{\rm eq}$, $M$ and $v_{\rm avg}$
We now focus on the radius evolution and its dependence on $T_{\rm eq},M,v_{\rm avg}$. After an extended exploration of parameters, here we show the main results for a representative set of models having $M=0.7,1,2,4\leavevmode\nobreak\ M_{j}$ and $T_{\rm eq}=1500,1750,2000,2250$ K, spanning the bulk of observed inflated HJs. For each combination of $M$ and $T_{\rm eq}$, there is a maximum value $v_{\rm avg}$ for which the runs are stable. We have looked for such values, and found an approximately linear dependence with $M$ and a much steeper one with $T_{\rm eq}$, which we have tentatively approximated (for the range of $M$ and $T_{\rm eq}$ here considered) as:
$v_{\rm avg}^{\rm stab}\sim 3000\frac{M_{j}}{M}\left(\frac{1500{\rm K}}{T_{\rm eq}}\right)^{6}\leavevmode\nobreak\ \frac{{\rm m}}{{\rm s}}\leavevmode\nobreak\ .$ (20)
@@
Note that this is not the inferred maximum value that fits the data and has not any a-priori justification: it is merely a functional form that approximately indicates the value of $v_{\rm avg}$ above which the code fails to converge. Runs with values $v_{\rm avg}\lesssim v_{\rm avg}^{\rm stab}$ show pronounced oscillations in e.g. the convective region thickness (see below), but the general trends of radius evolution as a function of the parameters, $R(t,M,T_{\rm eq},v_{\rm avg})$ are not affected by these details. The instability observed in extreme models is physical in nature and arises from a runaway process triggered by a certain excess of heat, as demonstrated, for example, in Batygin et al. (2011). In these cases, the planet undergoes expansion, eventually approaching Roche lobe overflow (a stage which is not followed up by our simulations).
@@
Each panel of Fig. 5 has shows models with a given $T_{\rm eq}=1750,2000,2250$ K (from left to right), and illustrates the radius evolution for various values of $M$ (indicated by different colours). and four values of $v_{\rm avg}$, equally spaced from zero to $v_{\rm avg}^{\rm stab}$, represented by increasingly darker shades. In each panel we show the cases which broadly encompass the observed radii around the value of $T_{\rm eq}$, indicated by the grey band. Each band delineates the observed mean radius $\bar{R}$ and its standard deviation $\sigma_{R}$, for the sample of HJs with a range of $T_{\rm eq}\pm 125$ K, as in Fig. 1. Here we mark it only for ages $t\geq$ 100 Myr for consistency with our sample selection (there are only a handful of younger HJs anyway). At the beginning of the evolution (after the relaxation phase mentioned at the end of Sect. 3.3) tracks start from a value close to the set initial radius, $R=4\leavevmode\nobreak\ R_{j}$ (see also App. C for general features of the cooling curves). For a given $M$ and $T_{\rm eq}$, different $v_{\rm avg}$ start showing difference when the Ohmic heating is turned on at 5 Myr.
@@
We note the strong dependence on the combined values of $T_{\rm eq}$ and $v_{\rm avg}$. The curves with the lightest shades (no Ohmic heating) show only slight inflation, which is larger for increasing $T_{\rm eq}$, due to the blanketing effect alone (see also benchmark models with no extra heating, App. C). For a given pair $(M,T_{\rm eq})$ there is a well-defined range of $v_{\rm avg}$ for which there is a noticeable effect on the radius. In our models, we manage to cover well the hottest HJs ($T_{\rm eq}\gtrsim 2000$ K) with values of $v_{\rm avg}$ of few hundred m/s maximum. If we consider lower values of $T_{\rm eq}$, we need instead to take the maximum allowed values of $\sim$ km/s, eq. (20), to cover the observed range of $R$, but models tend to fall short, compared to the more irradiated cases. Note also that, naturally, in our exploration of parameters, massive planets turn out being less inflated. This is related to the scaling laws and the role played by the dynamo field in the induction, as we discuss below.
@@
Figure 6: Evolution of radii (top row, with dots indicating the dynamo radius and the solid the planetary radius), estimated magnetic field produced in the dynamo region (second row, with dots for $B_{\rm dyn}$ and solid lines for the planetary surface, $B_{\rm atm}$, assuming $f_{\rm dip}=0.1$), Ohmic efficiency (third row), and $\sigma_{\rm atm}$ (bottom row). Each column represent a mass, from left to right: $M=0.7,1,4\leavevmode\nobreak\ M_{j}$. In all panels, colors indicate $T_{\rm eq}=1500,1750,2000,2250$ K (blue, green, orange, respectively), and four values $v_{\rm avg}=\{0,1,2,3\}/[(M/M_{j})(T_{\rm eq}/1500{\rm K})^{6}]$ km/s are indicated by increasingly darker shades. The curves of magnetic fields have been smoothen out by a Savitzky–Golay filter (window length 21, polynomial degree 3) to visually smooth out the frequent oscillations seen in the most heated/irradiated cases in light planets (see text).
@@
To understand more in depth these trends, Fig. 6 focuses on the most relevant physical quantities for the same set of simulations, considering $M=0.7,1,4\leavevmode\nobreak\ M_{j}$ (left to right columns), $T_{\rm eq}=1500,1750,2000,2250$ K (different colours), and the four same values $v_{\rm avg}$ mentioned above (different shades). In the top panels we show the nominal surface of the dynamo region in dashes, $R_{\rm dyn}=r(p_{\rm dyn})$, together with the planetary radii $R$ (solid lines). For all the cases, the absolute value of the dynamo radius is fairly constant in time and depends on the mass, due to the fact that the planet shrinking ($R$ decreasing) mostly affects the outer, less dense envelope. The difference between $R$ and $R_{\rm dyn}$ is more pronounced for internally hot planets, i.e. at early ages, with high $T_{\rm eq}$ and/or high Ohmic rate. Heavier planets (look at $4\leavevmode\nobreak\ M_{j}$ models, right column) show less difference due their higher gravity, with the dynamo radius being quite shallow (like in cold Jupiters) at Gyr ages.
@@
This trend has a consequence on the evolution of the dynamo and surface magnetic field, $B_{\rm dyn}$ and $B_{\rm atm}$, shown with dashed and solid lines, respectively, in the second row of Fig. 6. For a given value of $v_{\rm avg}$, $T_{\rm eq}$, the values of $B_{\rm dyn}$ increase with mass, due to the scaling laws. However, how this translates into a trend between mass and $B_{\rm atm}$ depends on the thickness of the outer envelope, i.e., how expanded are the layers outside the dynamo region, $R-R_{\rm dyn}$, which determines the difference between the dynamo and surface field, eq. (19). As a consequence, the surface field $B_{\rm atm}$ is always in the Jupiter range, $1-10$ G. Actually, larger deposited heat have the effect of decrease the surface field, due to the larger $R-R_{\rm dyn}$ and the effect of suppressing convection (causing also the stochastic oscillations in the $B$ curves), described below. Only massive planets, represented by the $4\leavevmode\nobreak\ M_{j}$ series (right column) have slightly higher values of $B_{\rm atm}\gtrsim 10$ G at Gyr ages. The trends of $B_{\rm dyn}$ and $B_{\rm atm}$ with mass differ with the studies by Yadav & Thorngren (2017); Thorngren (2024), probably because in that case they consider the Reiners et al. (2009) version of the scaling law, using the Ohmic luminosity as the one setting the convective heat flux (see a discussion in App. A).
@@
A very important consequence of the trend in the magnetic field is the evolution of the Ohmic efficiency, shown in the third row. In all models, it decreases in time, due to the $Q_{j}(t)\propto B_{\rm atm}^{2}(t)$ dependence, via the normalization of $J_{\rm atm}$, eq. (10). This feature is in contrast with classical Ohmic models, which assume a constant value of efficiency. As a consequence, our models do not predict the commonly found stalling of the evolution, i.e. the radius remaining more or less constant over the later stages in the evolution (e.g., Komacek & Youdin 2017), for which the volume-integrated heating balances the cooling. The effect is particularly pronounced for the massive cases, with a drop of $\sim 1.5$ orders of magnitude between $10^{8}$ and $10^{10}$ yr.
The typical Ohmic efficiency at Gyr as a function of $T_{\rm eq}$ preserves the same qualitative trend seen in previous statistical studies (Thorngren & Fortney, 2018; Sarkis et al., 2021), peaking at intermediate temperature (in our case, $T=1750$ K), although a more detailed study is needed for a quantitative comparison.

### v2

@@ 4.3 Estimates of dynamo magnetic fields
where $V$, $r_{i}$ and $r_{o}$ refer to the volume, inner and outer boundary of the convective region, $\mu_{0}$ is the magnetic permeability, $L={\rm min}(r_{o}-r_{i},H_{\rho})$, where $H_{\rho}$ is the density scale height, $H_{T}$ is the temperature scale height, $q_{c}(r)$ is the convective heat flux, and $f_{\rm ohm}$ is the fraction of Ohmic to total dissipation in the dynamo region (which Christensen & Aubert (2006) find numerically to be in the range $\sim 0.3$-$0.7$, depending on the simulation). We evaluate the integral taking $r_{i}$ as the bottom of the envelope (outside the inert core), and $r_{o}=R_{\rm dyn}$, i.e. the radius corresponding to a pressure $p_{\rm dyn}:=10^{6}$ bar, taken as a proxy for the surface of the deep high-conductivity region hosting the dynamo. We note that, strictly speaking, the transition to the metallic hydrogen region corresponds to a $\rho$-$T$ line in the phase diagram, so that $p_{\rm dyn}$ depends slightly on $T$ (or $\rho$). However, since the evaluation of the field is performed via an integral, Eq. (18), a slight change of the upper bound $r_{o}$ implies minor modifications, certainly smaller than the overall intrinsic uncertainties of the scaling law. We checked that, with this recipe, we recover reasonable values of $B_{\rm dyn}\sim 30-50$ G for a Jupiter-like model (negligible irradiation, $M=1\leavevmode\nobreak\ M_{\rm j}$), at $4.5$ Gyr-old. In our code, we use $f_{\rm ohm}\sim 0.5$, in the middle of the typical range $f_{\rm ohm}\sim 0.3-0.8$ of e.g. the geodynamo simulations of Christensen & Aubert (2006) and the anelastic ones by Gastine & Wicht (2021), and on the upper range of cold Jupiter simulations by Elias-López et al. (2025b). In any case, we note that the systematic uncertainties in the absolute values of $B_{\rm dyn}(t)$ (which we can overall quantify of a factor $\sim 2$) affect the inferred values of $v_{\rm avg}$, but do not affect the general trends of the magnetic field evolution. We refer the reader to App. A for a comparison of different versions of the scaling law (Reiners et al., 2009; Reiners & Christensen, 2010), and some words of caution about their applicability to the HJ scenario.
Moreover, we estimated the average value of the magnetic field intensity at the surface (close to where winds circulate), which we use in eq. (10) by considering that the weights $b_{l}(R_{\rm dyn})$ (see Sect. 4.2.2):
@@
where the last equation considers as a only free parameter the dipolarity at the dynamo surface, $f_{\rm dip}$, for simplicity (and given the much larger source of uncertainties in the scaling law). The set of anelastic dynamo simulations of Elias-López et al. (2025b) found that the dipolarity (calculated over the entire dynamo region, or at its outer surface) is seen to suddenly increase at a few gigayears, switching from a multipolar- to a dipolar-dominated dynamo (see also Yadav et al. 2013 for a discussion of these two branches). Here, for simplicity, we employed $f_{\rm dip}=0.1$, constant in time, in the range of found by Elias-López et al. (2025b). According to that study, neglecting its time variation might overestimate the surface field at early ages. Using these values of $f_{\rm ohm}$ and $f_{\rm dip}$ allows us to roughly recover Jupiter values of magnetic field at the surface.
@@ 4.4 Numerical details about the heating term
We gradually activate the Ohmic source term $\epsilon_{\rm j}$, eq. (6), with a linear increase in time starting from zero at 5 Myr, to have it fully operative at 20 Myr. The reasons to neglect the Ohmic term at earlier ages are: (i) we only consider HJs with $t>100$ Myr, discarding the few younger ones (see Sect. 2); (ii) planetary formation, migration, and interaction with the disc are still happening, and the Ohmic model does not account for them; (iii) the internal structure and the radius is still sensitive to the (poorly constrained) initial entropy (i.e. initial radius), something irrelevant after $\sim 100$ Myr; (iv) at tless-than-or-similar-$t\lesssim$ Myr, the different versions of the scaling law differ substantially between each other (see App. A), and eq. (18) easily gives unphysical Ohmic efficiencies due to the large heat flux (which in turn is affected by the choice of the initial entropy).
@@
Figure 5: Evolution of the radius, $R$, for planets with a given $T_{\rm eq}=1750,2000,2250$ K (from left to right), different masses ($M=0.7,1,2,4$ $M_{\rm j}$ in blue, green, orange, and purple, respectively), and different values of $v_{\rm avg}$, marked by increasingly darker shades: $v_{\rm avg}=\{0,1000,2000,3000\}/[(M/M_{\rm j})(T_{\rm eq}/1500{\rm K})^{6}]$ m/s. In each panel, the grey band indicates the range of observed radii for HJs with $T_{\rm eq}$ around (±$\pm 125$ K), which is the value assumed in the models, $\bar{R}\pm\sigma_{R}$, as shown by grey background boxes in Fig. 1.
@@
Due to the non-trivial feedback between the background field and the internal heat, the code can show instabilities (for both numerical and physical reasons, see below) for high level of irradiation and/or heat deposition, visible as perturbations in the long-term evolutionary curves, and possibly implying early crashes in the simulations. To reduce the numerical artefacts, we use a reduced time and grid resolution, typically 2x or 4x the default one (i.e. delta_time_coeff and mesh_delta_coeff 0.5 or 0.25 instead on 1). We also implement an exponential decrease of the heat source for $p\lesssim 1$ bar, which helps the convergence in the structure solution and does not affect the results in terms of inflated radii (the fraction of the heat deposited in these outermost thin layers is minimal).
@@
Additionally, the source, $\epsilon_{\rm j}$, at a given time, $t$, can actually be implemented as the average of the Ohmic heat, $<\frac{Q^{n}_{\rm j}}{\rho^{n}}>_{t_{\rm avg}}$, over an interval much smaller than the typical cooling timescales, which are comparable to the age itself (we checked that such a time average does not affect the overall long-term trend of the radius evolution). This simplified multistep-like method, adapted to our case, smooths out the oscillations in the instantaneous Ohmic heating rate, which are due to intermittent convection and dynamo field (see Sect. 5.2) and stabilise the code. In other words, we numerically mitigate the stiffness of the source term. We notice that in our specific case, the use of the implicit scheme for extra energy sources (with the recently implemented auto differentiation, Jermyn et al. 2023) does not improve the stability due to the intrinsically variable instantaneous heat source. Therefore, we opted for the more manual solution described above.
@@ 5.1 Dependence on $T_{\rm eq}$, $M$, and $v_{\rm avg}$
In this section, we focus on the radius evolution and its dependence on $T_{\rm eq},M,v_{\rm avg}$. After an extended exploration of parameters, we show the main results for a representative set of models with $M=0.7,1,2,4\leavevmode\nobreak\ M_{\rm j}$ and $T_{\rm eq}=1500,1750,2000,2250$ K spanning the bulk of the observed inflated HJs. For each combination of $M$ and $T_{\rm eq}$, there is a maximum value, $v_{\rm avg}$, for which the runs are stable. We looked for such values and found an approximately linear dependence with $M$ and a much steeper dependence with $T_{\rm eq}$, which we have tentatively approximated (for the range of $M$ and $T_{\rm eq}$ here considered) as
$v_{\rm avg}^{\rm stab}\sim 3000\frac{M_{\rm j}}{M}\left(\frac{1500\,{\rm K}}{T_{\rm eq}}\right)^{6}\leavevmode\nobreak\ \frac{{\rm m}}{{\rm s}}\leavevmode\nobreak\ .$ (20)
@@
We note that this is not the inferred maximum value that fits the data and that it does not have any a priori justification. It is merely a functional form that approximately indicates the value of $v_{\rm avg}$ above which the code fails to converge. Runs with values $v_{\rm avg}\lesssim v_{\rm avg}^{\rm stab}$ show pronounced oscillations in, for example, the convective region thickness (see below), but the general trends of radius evolution as a function of the parameters $R(t,M,T_{\rm eq},v_{\rm avg})$ are not affected by these details. The instability observed in extreme models is physical in nature and arises from a runaway process triggered by a certain excess of heat, as demonstrated, for example, in Batygin et al. (2011). In these cases, the planet undergoes expansion, eventually approaching Roche-lobe overflow (a stage which is not followed up by our simulations).
@@
Each panel of Fig. 5 has shows models with a given $T_{\rm eq}=1750,2000,2250$ K (from left to right), and illustrates the radius evolution for various values of $M$ (indicated by different colours). and four values of $v_{\rm avg}$, equally spaced from zero to $v_{\rm avg}^{\rm stab}$, represented by increasingly darker shades. In each panel we show the cases which broadly encompass the observed radii around the value of $T_{\rm eq}$, indicated by the grey band. Each band delineates the observed mean radius $\bar{R}$ and its standard deviation $\sigma_{R}$, for the sample of HJs with a range of $T_{\rm eq}\pm 125$ K, as in Fig. 1. Here we mark it only for ages $t\geq$ 100 Myr for consistency with our sample selection (there are only a handful of younger HJs anyway). At the beginning of the evolution (after the relaxation phase mentioned at the end of Sect. 3.3) tracks start from a value close to the set initial radius, $R=4\leavevmode\nobreak\ R_{\rm j}$ (see also App. C for general features of the cooling curves). For a given $M$ and $T_{\rm eq}$, different $v_{\rm avg}$ start showing difference when the Ohmic heating is turned on at 5 Myr.
@@
We note the strong dependence on the combined values of $T_{\rm eq}$ and $v_{\rm avg}$. The curves with the lightest shades (no Ohmic heating) show only slight inflation, which is larger for increasing $T_{\rm eq}$, due to the blanketing effect alone (see also benchmark models with no extra heating, App. C). For a given pair $(M,T_{\rm eq})$ there is a well-defined range of $v_{\rm avg}$ for which there is a noticeable effect on the radius. In our models, we manage to cover well the hottest HJs ($T_{\rm eq}\gtrsim 2000$ K) with values of $v_{\rm avg}$ of a few hundred metres per second maximum. If we consider lower values of $T_{\rm eq}$, we need instead to take the maximum allowed values of $\sim$ km/s, eq. (20), to cover the observed range of $R$, but models tend to fall short, compared to the more irradiated cases. We also note that, naturally, in our exploration of parameters, massive planets turn out being less inflated. This is related to the scaling laws and the role played by the dynamo field in the induction, as we discuss below.
@@
Figure 6: Evolution of radii (top row, with dots indicating the dynamo radius and the solid the planetary radius), estimated magnetic field produced in the dynamo region (second row, with dots for $B_{\rm dyn}$ and solid lines for the planetary surface, $B_{\rm atm}$, assuming $f_{\rm dip}=0.1$), Ohmic efficiency (third row), and $\sigma_{\rm atm}$ (bottom row). Each column represent a mass, from left to right: $M=0.7,1,4\leavevmode\nobreak\ M_{\rm j}$. In all panels, colours indicate $T_{\rm eq}=1500,1750,2000,2250$ K (blue, green, orange, respectively), and four values $v_{\rm avg}=\{0,1,2,3\}/[(M/M_{\rm j})(T_{\rm eq}/1500{\rm K})^{6}]$ km/s are indicated by increasingly darker shades. The curves of magnetic fields have been smoothed out by a Savitzky–Golay filter (window length 21, polynomial degree 3) to visually smooth out the frequent oscillations seen in the most heated and irradiated cases in light planets (see text).
@@
To understand more in depth these trends, Fig. 6 focuses on the most relevant physical quantities for the same set of simulations, considering $M=0.7,1,4\leavevmode\nobreak\ M_{\rm j}$ (left to right columns), $T_{\rm eq}=1500,1750,2000,2250$ K (different colours), and the four same values $v_{\rm avg}$ mentioned above (different shades). In the top panels we show the nominal surface of the dynamo region in dashes, $R_{\rm dyn}=r(p_{\rm dyn})$, together with the planetary radii $R$ (solid lines). For all the cases, the absolute value of the dynamo radius is fairly constant in time and depends on the mass, due to the fact that the planet shrinking ($R$ decreasing) mostly affects the outer, less dense envelope. The difference between $R$ and $R_{\rm dyn}$ is more pronounced for internally hot planets, i.e. at early ages, with high $T_{\rm eq}$ and/or high Ohmic rate. Heavier planets (e.g. the $4\leavevmode\nobreak\ M_{\rm j}$ models, right column) show fewer differences due their higher gravity, with the dynamo radius being quite shallow (as in cold Jupiters) at gigayear ages.
@@
This trend has a consequence on the evolution of the dynamo and surface magnetic field, $B_{\rm dyn}$ and $B_{\rm atm}$, shown with dashed and solid lines, respectively, in the second row of Fig. 6. For a given value of $v_{\rm avg}$, $T_{\rm eq}$, the values of $B_{\rm dyn}$ increase with mass, due to the scaling laws. However, how this translates into a trend between mass and $B_{\rm atm}$ depends on the thickness of the outer envelope, i.e. how expanded are the layers outside the dynamo region, $R-R_{\rm dyn}$, which determines the difference between the dynamo and surface field, eq. (19). As a consequence, the surface field $B_{\rm atm}$ is always in the Jupiter range, $1-10$ G. Actually, larger deposited heat have the effect of decrease the surface field, due to the larger $R-R_{\rm dyn}$ and the effect of suppressing convection (causing also the stochastic oscillations in the $B$ curves), described below. Only massive planets, represented by the $4\leavevmode\nobreak\ M_{\rm j}$ series (right column) have slightly higher values of $B_{\rm atm}\gtrsim 10$ G at gigayear ages. The trends of $B_{\rm dyn}$ and $B_{\rm atm}$ with mass differ with the studies by Yadav & Thorngren (2017); Thorngren (2024), probably because in that case they consider the Reiners et al. (2009) version of the scaling law, using the Ohmic luminosity as the one setting the convective heat flux (see a discussion in App. A).
@@
A very important consequence of the trend in the magnetic field is the evolution of the Ohmic efficiency, shown in the third row. In all models, it decreases in time, due to the $Q_{\rm j}(t)\propto B_{\rm atm}^{2}(t)$ dependence, via the normalisation of $J_{\rm atm}$, eq. (10). This feature is in contrast with classical Ohmic models, which assume a constant value of efficiency. As a consequence, our models do not predict the commonly found stalling of the evolution, i.e. the radius remaining more or less constant over the later stages in the evolution (e.g. Komacek & Youdin 2017), for which the volume-integrated heating balances the cooling. The effect is particularly pronounced for the massive cases, with a drop of $\sim 1.5$ orders of magnitude between $10^{8}$ and $10^{10}$ yr.
The typical Ohmic efficiency at gigayears as a function of $T_{\rm eq}$ preserves the same qualitative trend seen in previous statistical studies (Thorngren & Fortney, 2018; Sarkis et al., 2021), peaking at intermediate temperature (in our case, $T=1750$ K), although a more detailed study is needed for a quantitative comparison.